OpenStudy (anonymous):

(c^-3)(d^-2)/(c-4)(d-5) = (c^-3)/(c^-4) x (d^-2)/(d^-5) Where do i go from here?

OpenStudy (anonymous):

\[\frac{\frac{1}{c^{3}d^{2}}}{(c-4)(c-5)}\]is this your question?

OpenStudy (anonymous):

\[C(-3) \times D(-5) \over C(-4) \times D(-5)\]

OpenStudy (anonymous):

thats the original question.

jhonyy9 (jhonyy9):

- so i think not is this ,i think there are missed the exponents signs ,i think that -4 and -5 are exponents too

jhonyy9 (jhonyy9):

(c^-3)(d^-2) 1/((c^3)(d^2)) 1 ((c^4)(d^5)) ------------ = --------------- = ----------- * ------------ = c^-4)(d^-5) 1/((c^4)(d^5)) ((c^3)(d^2)) 1 = c^(4-3)*d^(5-2) = c*d^3

OpenStudy (anonymous):

the equation i wroe is exctly how it is in the book. the -4n -5 are not exponents.